# When predict can also explain: few-shot prediction to select better neural latents

### v3

@@ Abstract
Latent variable models serve as powerful tools to infer underlying dynamics from observed neural activity. Ideally, the inferred dynamics should align with true ones. However, due to the absence of ground truth data, prediction benchmarks are often employed as proxies. One widely-used method, co-smoothing, involves jointly estimating latent variables and predicting observations along held-out channels to assess model performance. In this study, we reveal the limitations of the co-smoothing prediction framework and propose a remedy. In a student-teacher setup with Hidden Markov Models, we demonstrate that the high co-smoothing model space encompasses models with arbitrary extraneous dynamics in their latent representations. To address this, we introduce a secondary metric — few-shot co-smoothing, performing regression from the latent variables to held-out channels in the data using fewer trials. Our results indicate that among models with near-optimal co-smoothing, those with extraneous dynamics underperform in the few-shot co-smoothing compared to ’minimal’ models that are devoid of such dynamics. We provide analytical insights into the origin of this phenomenon and further validate our findings on real neural data using two state-of-the-art methods: LFADS and STNDT. In the absence of ground truth, we suggest a novel measure to validate our approach. By cross-decoding the latent variables of all model pairs with high co-smoothing, we identify models with minimal extraneous dynamics. We find a correlation between few-shot co-smoothing performance and this new measure. In summary, we present a novel prediction metric designed to yield latent variables that more accurately reflect the ground truth, offering a significant improvement for latent dynamics inference. Code available here.
Machine Learning, ICML
1 Introduction
@@
In neuroscience, we often have access to simultaneously recorded neurons during certain behaviors. These observations, denoted ${\bm{X}}$, offer a window onto the actual hidden (or latent) dynamics of the relevant brain circuit, denoted ${\bm{Z}}$ (Vyas et al., 2020). Although, in general, these dynamics can be complex and high-dimensional, capturing them in a concrete mathematical model opens doors to reverse-engineering, revealing simpler explanations and insights (Barak, 2017; Sussillo & Barak, 2013). Inferring a model of the ${\bm{Z}}$ variables, also known as latent variable modeling (LVM), is part of the larger field of system identification with applications in many areas outside of neuroscience, such as fluid dynamics (Vinuesa & Brunton, 2022) and finance (Bauwens & Veredas, 2004).
@@
Because we don’t have ground truth for ${\bm{Z}}$, prediction metrics on held-out parts of $x$ are commonly used as a proxy (Pei et al., 2021). However, it has been noted that prediction and explanation are often distinct endeavors (Shmueli, 2010). For instance, (Versteeg et al., 2023) use an example where ground truth is available to show how different models that all achieve good prediction nevertheless have varied latents that can differ from the ground truth. Such behavior might be expected when using highly expressive models with large latent spaces. Bad prediction with good latents is demonstrated by (Koppe et al., 2019) for the case of chaotic dynamics.
@@
Various regularisation methods on the latents have been suggested to improve the similarity of ${\bm{Z}}$ to the ground truth, such as recurrence and priors on external inputs (Pandarinath et al., 2018), low-dimensionality of trajectories (Sedler et al., 2022), low-rank connectivity (Valente et al., 2022; Pals et al., 2024), injectivity constraints from latent to predictions (Versteeg et al., 2023), low-tangling (Perkins et al., 2023), and piecewise-linear dynamics (Koppe et al., 2019). However, the field lacks a quantitative, prediction-based metric that credits the simplicity of the latent representation—an aspect essential for interpretability and ultimately scientific discovery, while still enabling comparisons across a wide range of LVM architectures.
@@
Here, we characterize the diversity of model latents achieving high co-smoothing, a standard prediction-based framework for Neural LVMs, and demonstrate potential pitfalls of this framework. We propose a few-shot variant of co-smoothing which, when used in conjunction with co-smoothing, differentiates varying latents. We verify this approach both on synthetic toy problems and state-of-the-art methods on neural data, providing an analytical explanation of why it works in a simple setting.
2 Related Work
Our work builds on recent developments in Neural LVMs for the discovery of latent structure in noisy neural data on single trials. We refer the reader to (Pei et al., 2021) supplementary table 3 for a comprehensive list of Neural LVMs published from 2008-2021. Central to our work is the co-smoothing procedure, which evaluates models based on the prediction of activity from held-out neurons provided held-in neuron activity from the same trial. Co-smoothing was first introduced in (Yu et al., 2008) and (Macke et al., 2011) for the validation of GPFA as a Neural LVM.
(Pei et al., 2021) curated four datasets of neural activity recorded from behaving monkeys and established a framework to evaluate co-smoothing among other prediction-based metrics on several models in the form of a standardized benchmark and competition.
In contrast to prediction approaches, a parallel line of work focuses on explaining and validating Neural LVMs on synthetic data, enabling direct comparison with the ground truth (Sedler et al., 2022; Brenner et al., 2022; Durstewitz et al., 2023). (Versteeg et al., 2023) validated their method with both ground truth and neural data, demonstrating high predictive performance with low-dimensional latents.
A concept we introduce is cross-decoding across a population of models to find the most parsimonious representation. Several works compare representations of large model populations (Maheswaranathan et al., 2019; Morcos et al., 2018). They apply Canonical Correlation Analysis (CCA), a symmetric measure of representational similarity, whereas we use regression, which is not symmetric. The application to Neural LVMs may be novel.
One related approach comparing representations in deep neural networks is stitching components of separately trained (and subsequently frozen) models into a composite model using a simple linear layer (Lenc & Vedaldi, 2015; Bansal et al., 2021).
Central to our work is the concept of few-shot learning a decoder from a frozen intermediate representation. (Sorscher et al., 2022) developed a theory of geometric properties of representations that enables few-shot generalization to novel classes. They identified the geometric properties that determine a signal-to-noise ratio for classification, which dictates few-shot performance. While this setting differs from ours, links between our works are a topic for future research. To our knowledge, the use of few-shot generalization as a means to identify interpretable latent representations, particularly for Neural LVMs, is a novel idea.
3 Co-smoothing: a cross-validation framework
@@
Let ${\bm{X}}\in{\mathbb{Z}}_{\geq 0}^{T\times N}$ be spiking neural activity of $N$ channels recorded over a finite window of time, i.e., a trial, and subsequently quantised into $T$ time-bins. ${X}_{t,n}$ represents the number of spikes in channel $n$ during time-bin $t$. The dataset $\mathcal{X}:=\{{\bm{X}}^{(i)}\}_{i=1}^{S}$, partitioned as $\mathcal{X}^{\text{train}}$ and $\mathcal{X}^{\text{test}}$, consists of $S$ trials of the experiment. The latent-variable model (LVM) approach posits that each time-point in the data ${\bm{X}}^{(i)}_{t,:}$ is a noisy measurement of a latent state ${\bm{Z}}^{(i)}_{t,:}$.
@@
To infer the latent trajectory ${\bm{Z}}$ is to learn a mapping $f:{\bm{X}}\mapsto{\bm{Z}}$. On what basis do we validate the inferred ${\bm{Z}}$? We have no ground truth on ${\bm{Z}}$, so instead we test the ability of ${\bm{Z}}$ to predict unseen or held-out data. Data may be held-out in time, e.g., predicting future data points from the past, or in space, e.g., predicting neural activities of one set of neurons (or channels) based on those of another set. The latter is called co-smoothing (Pei et al., 2021).
@@
${\bm{Z}}_{t,:}$ $=f({\bm{X}}_{:,\text{in}};t)$ (1)
$R_{t,\text{out}}$ $=g({\bm{Z}}_{t,:}),$ (2)
where the subscripts ‘in’ and ‘out’ denote partitions of the neurons. During evaluation, the held-out data from test trials ${\bm{X}}_{:,\text{out}}$ is compared to the rate-predictions ${\bm{R}}_{:,\text{out}}$ from the model using the co-smoothing metric $\mathcal{Q}$ defined as the normalised log-likelihood, given by:
@@
where $\mathcal{L}$ is poisson log-likelihood, $\bar{r}_{n}=\frac{1}{TS}\sum_{i}\sum_{t}{X}^{(i)}_{t,n}$ is a the mean rate for channel $n$, and $\mu_{n}:=\sum_{i}\sum_{t}{X}^{(i)}_{t,n}$ is the total number of spikes, following (Pei et al., 2021).
@@
using $\mathcal{X}^{\text{train}}$, without access to the test trials from $\mathcal{X}^{\text{test}}$. For claritry, apart from equation 5, we report only $\mathcal{Q}^{\text{test}}$, omitting the superscript.

### v4

@@ Abstract
Latent variable models serve as powerful tools to infer underlying dynamics from observed neural activity. Ideally, the inferred dynamics should align with true ones. However, due to the absence of ground truth data, prediction benchmarks are often employed as proxies. One widely-used method, co-smoothing, involves jointly estimating latent variables and predicting observations along held-out channels to assess model performance. In this study, we reveal the limitations of the co-smoothing prediction framework and propose a remedy. Using a student-teacher setup, we demonstrate that models with high co-smoothing can have arbitrary extraneous dynamics in their latent representations. To address this, we introduce a secondary metric — few-shot co-smoothing, performing regression from the latent variables to held-out neurons in the data using fewer trials. Our results indicate that among models with near-optimal co-smoothing, those with extraneous dynamics underperform in the few-shot co-smoothing compared to ‘minimal’ models that are devoid of such dynamics. We provide analytical insights into the origin of this phenomenon and further validate our findings on four standard neural datasets using a state-of-the-art method: STNDT. In the absence of ground truth, we suggest a novel measure to validate our approach. By cross-decoding the latent variables of all model pairs with high co-smoothing, we identify models with minimal extraneous dynamics. We find a correlation between few-shot co-smoothing performance and this new measure. In summary, we present a novel prediction metric designed to yield latent variables that more accurately reflect the ground truth, offering a significant improvement for latent dynamics inference.
Author summary
The availability of large scale neural recordings encourages the development of methods to fit models to data. How do we know that the fitted models are loyal to the true underlying dynamics of the brain? A common approach is to use prediction scores that use one part of the available data to predict another part. The advantage of predictive scores is that they are general: a wide variety of modelling methods can be evaluated and compared against each other. But does a good predictive score guarantee that we capture the true dynamics in the model?
We investigate this by generating synthetic neural data from one model, fitting another model to it, ensuring a high predictive score, and then checking if the two are similar. The result: only partially. We find that the high scoring models always contain the truth, but may also contain additional ‘made-up’ features. We remedy this issue with a secondary score that tests the model’s generalisation to another set of neurons with just a few examples. We demonstrate its applicability with synthetic and real neural data.
Introduction
@@
In neuroscience, we often have access to simultaneously recorded neurons during certain behaviors. These observations, denoted $\bm{X}$}, offer a window onto the actual hidden (or latent) dynamics of the relevant brain circuit, denoted $\bm{Z}$} [1]. Although, in general, these dynamics can be complex and high-dimensional, capturing them in a concrete mathematical model opens doors to reverse-engineering, revealing simpler explanations and insights [2, 3]. Inferring a model of the $\bm{Z}$} variables, $\hat{\bm{Z}}$, also known as latent variable modeling (LVM), is part of the larger field of system identification with applications in many areas outside of neuroscience, such as fluid dynamics [4] and finance [5].
@@
Because we don’t have ground truth for $\bm{Z}$}, prediction metrics on held-out parts of $\bm{X}$} are commonly used as a proxy [6]. However, it has been noted that prediction and explanation are often distinct endeavors [7]. For instance, [8] use an example where ground truth is available to show how different models that all achieve good prediction nevertheless have varied latents that can differ from the ground truth. Such behavior might be expected when using highly expressive models with large latent spaces. Bad prediction with good latents is demonstrated by [9] for the case of chaotic dynamics.
@@
Various regularisation methods on the latents have been suggested to improve the similarity of $\bm{Z}$} to the ground truth, such as recurrence and priors on external inputs [10], low-dimensionality of trajectories [11], low-rank connectivity [12, 13], injectivity constraints from latent to predictions [8], low-tangling [14], and piecewise-linear dynamics [15]. However, the field lacks a quantitative, prediction-based metric that credits the simplicity of the latent representation—an aspect essential for interpretability and ultimately scientific discovery, while still enabling comparisons across a wide range of LVM architectures.
@@
Here, we characterize the diversity of model latents achieving high co-smoothing, a standard prediction-based framework for Neural LVMs, and demonstrate potential pitfalls of this framework (see Methods for a glossary of terms). We propose a few-shot variant of co-smoothing which, when used in conjunction with co-smoothing, differentiates varying latents. We verify this approach both on synthetic data settings and a state-of-the-art method on neural data, providing an analytical explanation of why it works in simple settings.
Co-smoothing: a cross-validation framework
@@
Let ${\bm{X}}\in{\mathbb{Z}}_{\geq 0}^{T\times N}$ be spiking neural activity of $N$ channels recorded over a finite window of time, i.e., a trial, and subsequently quantised into $T$ time-bins. ${X}_{t,n}$ represents the number of spikes in channel $n$ during time-bin $t$. The dataset $\mathcal{X}:=\{{\bm{X}}^{(i)}\}_{i=1}^{S}$, partitioned as $\mathcal{X}^{\text{train}}$ and $\mathcal{X}^{\text{test}}$, consists of $S$ trials of the experiment. The latent-variable model (LVM) approach posits that each time-point in the data $\bm{x}^{(i)}_{t}$ is a noisy measurement of a latent state $\bm{z}^{(i)}_{t}$.
@@
To infer the latent trajectory $\bm{Z}$} is to learn a mapping $f:{\bm{X}}\mapsto\hat{\bm{Z}}$. On what basis do we validate the inferred $\hat{\bm{Z}}$? We cannot access the ground truth $\bm{Z}$}, so instead we test the ability of $\hat{\bm{Z}}$ to predict unseen or held-out data. Data may be held-out in time, e.g., predicting future data points from the past, or in space, e.g., predicting neural activities of one set of neurons (or channels) based on those of another set. The latter is called co-smoothing [6].
@@
$\bm{\hat{z}}_{t}$ $=f({\bm{X}}_{:,\text{in}};t)$ (1)
$\bm{r}_{t,\text{out}}$ $=g(\bm{\hat{z}}_{t}),$ (2)
where the subscripts ‘in’ and ‘out’ denote partitions of the neurons (Fig. 1B). During evaluation, the held-out data from test trials ${\bm{X}}_{:,\text{out}}$ is compared to the rate-predictions ${\bm{R}}_{:,\text{out}}$ from the model using the co-smoothing metric $\mathcal{Q}$ defined as the normalised log-likelihood, given by:
@@
where $\mathcal{L}$ is poisson log-likelihood, $\bar{r}_{n}=\frac{1}{TS}\sum_{i}\sum_{t}{X}^{(i)}_{t,n}$ is a the mean rate for channel $n$, and $\mu_{n}:=\sum_{i}\sum_{t}{X}^{(i)}_{t,n}$ is the total number of spikes, following [6].
@@
using $\mathcal{X}^{\text{train}}$, without access to the test trials from $\mathcal{X}^{\text{test}}$. For clarity, apart from equation 5, we report only $\mathcal{Q}^{\text{test}}$, omitting the superscript.
Fig 1: Prediction framework and its relation to ground truth. A. To evaluate a neural LVM with co-smoothing, the dataset is partitioned along the neurons and trials axes. B. The held-in neurons are used to infer latents $\hat{z}$, while the held-out serve as targets for evaluation. The encoder $f$ and decoder $g$ are trained jointly to maximise co-smoothing $\mathcal{Q}$. After training, the composite mapping $g\circ f$ is evaluated on the test set. C. We hypothesise that models with high co-smoothing may have an asymmetric relationship to the true system, ensuring that model representation contains the ground truth, but not vice-versa. We reveal this in a synthetic student(S)-teacher(T) setting by the unequal performance of regression on the states in the two directions. $\mathcal{D}_{u\rightarrow v}$ denote decoding error of model $v$ latents $\bm{z}_{v}$ from model $u$ latents $\bm{z}_{u}$.
Good co-smoothing does not guarantee correct latents
It is common to assume that being able to predict held-out parts of $\bm{X}$} will guarantee that the inferred latent aligns with the true one [16, 6, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 14]. To test this assumption, we use a student-teacher scenario where we know the ground truth. To compare how two models ($u,v$) align, we infer the latents of both from $\mathcal{X}^{\text{test}}$, then do a regression from latents of $u$ to $v$. The regression error is denoted $\mathcal{D}_{u\rightarrow v}$ (i.e. $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ for teacher to student decoding). Contrary to the above assumption, we hypothesize that good prediction guarantees that the true latents are contained within the inferred ones (low $\mathcal{D}_{\text{S}\rightarrow\text{T}}$), but not vice versa (Fig. 1C). It is possible that the inferred latents possess additional features, unexplained by the true latents (high $\mathcal{D}_{\text{T}\rightarrow\text{S}}$).
We demonstrate this phenomenon in three different student-teacher scenarios: task-trained RNNs, Hidden Markov Models (HMMs) and linear gaussian state space models. We start with RNNs, as they are a standard tool to investigate computation through dynamics in neuroscience [29], and expand upon the other models in the appendix. A $128$-unit RNN teacher (Methods) is trained on a 2-bit flip-flop task, inspired by working memory experiments. The network receives input pulses and has to maintain the identity of the last pulse (see Methods). The student is a sequential autoencoder, where the encoder $f$ is composed of a neural network that converts observations into an initial latent state, and another recurrent neural network that advances the latent state dynamics [29] (see Methods).
We generated a dataset of observations from this teacher, and then trained $30$ students with latent-dimensionality 3643-64 on the same teacher data using gradient-based methods (see Methods). Co-smoothing scores of students increased with the size of the latents, but are high for models in the range of 5-15 dimensional latents (S1 Fig.). Consistent with our hypothesis, the ability to decode the teacher from the student was highly correlated to the co-smoothing score (Fig. 2 top left). In contrast, the ability to decode the student from the teacher has a very different pattern. For students with low co-smoothing, this decoding is good – but meaningless. For students with high co-smoothing, there is a large variability, and little correlation to the co-smoothing score (Fig. 2 top right). In this simple example, it would seem that one only needs to increase the dimensionality of the latent until co-smoothing saturates. This minimal value would satisfy both demands. This is not the case for real data, as will be shown below.
What is it about a student model, that produces good co-smoothing with the wrong latents? It’s easiest to see this in a setting with discrete latents, so we first show the HMM teacher and two exemplar students – named “Good” and “Bad” (marked by green and red arrows in S3 Fig.AB) – and visualise their states and transitions using graphs in Fig. 2. The teacher is a cycle of 4 steps. The good student contains such a cycle (orange), and the initial distribution is restricted to that cycle, rendering the other states irrelevant. In contrast, the bad student also contains this cycle (orange), but the initial distribution is not consistent with the cycle, leading to an extraneous branch converging to the cycle, as well as a departure from the main cycle (both components in dark colour). Note that this does not interfere with co-smoothing, because the emission probabilities of the extra states are consistent with true states, i.e., the emission matrix conceals the extraneous dynamics. In the RNN, we see a qualitatively similar picture, with the bad students having dynamics in task-irrelevant dimensions (Fig. 2 “Bad” S).